40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 24 11:04:53 2022, JC, Update, VAC, Beginning Pumpdown PUMPDOWN-2022-05-24_16-57-59.pngC1VAC_Screenshot_2022-05-24_16-59-27.png
    Reply  Wed May 25 16:38:27 2022, yuta, Update, SUS, Suspensions quick health check Screenshot_2022-05-25_16-46-13.pngOSEMcheck.JPG
       Reply  Tue Jun 7 17:26:21 2022, yuta, Update, SUS, ITMY ULCOIL mystery not solved ULCoilInductance.JPGULCoilResistance.JPGWrongSetting.JPG
          Reply  Tue Jun 7 19:13:12 2022, yuta, Update, SUS, Agreement on suspension damping loop polarity conventions NoteBeforeChanges.JPG
             Reply  Fri Jun 10 15:02:04 2022, yuta, Update, SUS, Servo gain sign flipped for MC1 WFS relief Screenshot_2022-06-10_15-04-46.png
          Reply  Tue Jun 7 19:40:45 2022, Anchal, Update, SUS, ITMY changed output matrix to disable use of UL coil ITMY_PIT_to_POS-PIT-YAW_Coupling.pdfITMY_YAW_to_POS-PIT-YAW_Coupling.pdf
          Reply  Wed Jun 8 18:16:20 2022, yuta, Update, SUS, ITMY ULCOIL mystery: Coil driver swap test Screenshot_2022-06-08_18-10-44.pngSWAP.JPG
             Reply  Thu Jun 9 23:08:39 2022, rana, Update, SUS, ITMY ULCOIL mystery: Coil driver swap test 
                Reply  Fri Jun 10 13:02:14 2022, yuta, Update, SUS, ITMY ULCOIL mystery: Coil driver swap test 
                   Reply  Fri Jun 10 15:04:23 2022, rana, Update, SUS, ITMY ULCOIL mystery: Coil driver swap test 
                      Reply  Fri Jun 10 21:10:01 2022, yuta, Update, SUS, ITMY ULCOIL mystery: Coil driver swap test 
                         Reply  Wed Jun 15 17:03:17 2022, yuta, Update, SUS, ITMY ULCOIL issue solved, loose connection in sat amp box Measured16Ohms.JPGSatAmpTerminator.JPGBAD.JPGSUCCESSFUL_KICK.png
       Reply  Wed Jun 15 15:45:37 2022, yuta, Update, SUS, LO1 LLCOIL now working, it was loose connection Measurement.JPGBAD.JPGScreenshot_2022-06-15_15-59-05.png
    Reply  Wed Jun 15 07:26:35 2022, JC, Update, VAC, Beginning Pumpdown 
Message ID: 16873     Entry time: Wed May 25 16:38:27 2022     In reply to: 16871     Reply to this: 16896   16919
Author: yuta 
Type: Update 
Category: SUS 
Subject: Suspensions quick health check 

[JC, Yuta]

We did a quick health check of suspesions after the pump down.

Summary:
 - ITMX LRSEN is too bright (~761) and not responding to any optic motions (we knew this before the pump down)
 - ITMY ULCOIL is not working
 - LO1 LLCOIL is not working
 - Damping loops need to be retuned, especially for ETMY (too much damping), SRM, PR3 and AS4 (damping too weak)
 - MC1 sensor outputs are minus instead of plus
 - LO2 OSEMs got stuck during the pump down, but now it is free after some kicks. OSEM sensorr values almost came back (see attached)

What we did:
 1. Kicked optics with C1:SUS-{optic}_{UL,LL,UR,LR,SD}COIL_OFFSET one by one with offsets of +/- 10000 (or 100000), and checked if C1:SUS-{optic}_{UL,LL,UR,LR,SD}SEN_OUT16 move in both directions.

 2. Check if the optic damps nicely.

 3. Attached photo of the note is the result.

Attachment 1: Screenshot_2022-05-25_16-46-13.png  47 kB  Uploaded Wed May 25 17:47:23 2022  | Hide | Hide all
Screenshot_2022-05-25_16-46-13.png
Attachment 2: OSEMcheck.JPG  3.990 MB  Uploaded Wed May 25 20:26:28 2022  | Hide | Hide all
OSEMcheck.JPG
ELOG V3.1.3-